OpenStudy (anonymous):

If point C is the midpoint of AB, what is the coordinate of point B? (3, 1) (4, 2) (4, 4) (3, 0)

OpenStudy (anonymous):

Here is the picture

OpenStudy (e.mccormick):

Well, if it is the midpoint, then it is half way up and half way over. So, you find the total change in x and y, then take half of that, and add it to the bottom point. That is one way to do it. Or you can add the points and halv that.

OpenStudy (anonymous):

Thank you both so much.

OpenStudy (e.mccormick):

Ah, this is going from the midpoint, I was talking about finding the idpoint. My mistake. So you woul double the up over in this case.

OpenStudy (anonymous):

Let B=(x,y) Then 3 = (x+2)/2 0 = (y-4)/2

OpenStudy (anonymous):

Ok, so 4,4

OpenStudy (e.mccormick):

Yes. Sorry bout the mistake. Hope you get the idea they are related. Lets you answer any of them easily.

OpenStudy (anonymous):

Yes thank you

OpenStudy (amistre64):

my idea is: since midpoint is average of the other 2 points ..\[\frac{A+B}{2}=C\] \[B=2C-A\]

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

2C = 2(3,0) = (6 , 0) -A = -(2, -4) = (-2, 4) ---------- 2C-A = ( 4 , 4)

OpenStudy (anonymous):

Thank you, makes sense now